A population consists of the four measurements as follow:

5 9 10 8

What is the mean and standard deviation of the population?


1
Expert's answer
2022-05-11T08:58:47-0400

We have population values 5,9,10, 8, population size N=4.

Mean of population "(\mu)" = "\dfrac{5+9+10+8}{4}=8"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{4}(9+1+4+0)"


"=3.5"

"\sigma=\sqrt{\sigma^2}=\sqrt{3.5}\approx1.8708"